# Microscopic Mechanism of Anyon Superconductivity Emerging from Fractional Chern Insulators

### v1

@@ IV Microscopic lattice model
Next, we study the spin and charge correlation lengths when tuning across the critical point. We find the density correlation length, corresponding to the charge sector $(Q,S_{z})=(0,0)$, to diverge at the critical point as a function of the bond dimension. In contrast, the spin correlation length, $(Q,S_{z})=(0,\hbar)$, remains finite and quickly converges with bond dimension; see Fig 7(b). This suggests a finite spin gap across the transition, consistent with the continuous transition discussed in Sec. II. Furthermore, the ratio between the charge $(Q,S_{z})=(2,0)$ and charge $(Q,S_{z})=(1,\hbar/2)$ correlation length peaks at the critical point, growing with bond dimension and are, in particular, larger than the trivial expectation of 1/2, which arises when the charge-2 correlations would be only inherited from charge-1 correlations; Fig. 7(c). This behavior is expected when charge-2 excitations become energetically favorable compared to spinful charge-1 excitations.
@@ V Experimental implications
Therefore, from a coarse-grained point of view, both our setting and MoTe2 host the fractional $2/3$ state, a proximate superconductor on doping, and nearby insulators without Hall conductance. However, some important differences between these systems should be noted, even at the level of symmetries. We have assumed spin rotation symmetry, while the valley polarized MoTe2 has stongly-broken spin rotation symmetry due to spin valley locking. Can the two be related? The first observation is that since our model has a spin gap, weak breaking of spin-rotation symmetry leaves the phases intact. For instance, the topological order of the chiral spin liquid is stable when spin rotation symmetry is broken [29], despite the fact that spin transport is not quantized any longer. Next, we note that for the strong spin-rotation symmetry breaking relevant to tMoTe2, our mechanism provides a general guideline: if the 2/3 FCI is in proximity to a specific charge density wave - the semion crystal - then the charge 2e/3 anyon becomes energetically cheaper than a pair of separated e/3 quasiparticles. Doping such an FCI should favor superconductivity. Repulsive interactions are indeed expected to drive FCIs towards CDW order, the key is to realize the correct type of CDW. A promising strategy is to search the phase diagram of MoTe2 for insulating states that continue to break time reversal symmetry but exhibit a vanishing Hall response, which can be a potential candidate for this exotic CDW and examine if we have a continuous (or weakly first order) transition into the 2/3 FCI. Indeed, it has already been speculated that precisely such exotic CDWs could arise in MoTe2 [28]. Our framework links that putative CDW directly to enhanced Cooper pairing in the neighboring FCI, providing a concrete route to accounting for superconductivity.
@@
As an alternative scenario, superconductivity in this system could instead evolve from the intervening chiral metal. To distinguish these scenarios, it will be helpful to have more information on the energy ordering of quasiparticles. A simple (but indirect and imperfect) way to do this is via the observation (or absence) of daughter FCI plateaus, as the hierarchical sequence changes depending on which anyon is lowest energy. If the minimal charge 1/3 anyon is the cheapest, then one obtains the usual Haldane-Halperin hierarchy [59], and several of these are observed for $\nu<2/3$. This is consistent with that being the lowest charge excitation on the electron doping (smaller $\nu$) side. However on the opposite side, incompressible states at fillings predicted by the 2/3 anyon hierarchy, such as $\nu=4/5$ would provide strong evidence for 2/3 anyons to be cheaper for hole doping (see Appendix B), regardless of the mechanism. This energy ordering would be consistent with the observation of superconductivity for $\nu>2/3$. Other potential signatures of the energy ordering of anyons could come from spectroscopic probes [60, 61, 62, 63, 64, 65] or by measuring the electric charge via shot-noise measurements [66, 67] or via single electron transitor measurements [68, 69].
@@
In the MoTe2 experiments, a re-entrant integer QAH plateau was reported [58] between the 2/3 and 3/5 FQAH states. Although additional confirmation is warranted—particularly since edge-equilibration effects could influence the measurement—taking the observation at face value, this state could potentially arise from the usual plateau transition driven by 1/3 anyons. In our formalism, stacking a Chern-1 QAH layer on top of the 2/5 to semion-crystal transition also reproduces such a feature (see Appendix C).
@@ VI Outlook & Discussions
One of the biggest surprises of our work is the numerical discovery of the semion crystal in simple Hubbard models at fractional filling. One may view this phase as a period-three electron CDW on which the residual spins form a chiral spin liquid, akin to the cluster-Mott physics proposed for 1T-TaS2 [71]. Elucidating the properties and stability of such semion crystals in other fractional fillings and lattice geometries is an interesting direction.
@@
Acknowledgments.— We thank J. Dong, P.J. Ledwith, J. Feldmeier, G. Shavit, J. May-Mann, T. Helbig, and J. Alicea for insightful discussions. A.V. would like to thank Mike Zaletel, Stefan Divic, Zhaoyu Han and Yahui Zhang for discussions and collaborations on related topics. C.K. acknowledges funding from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 217884). A.V. is supported by a Simons Investigator award, the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651440, A.V.). F.P. and M.K. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, TRR 360 – 492547816 and DFG grants No. KN1254/1-2, KN1254/2-1, the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 851161), the European Union (grant agreement No 101169765), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@
Appendix A Charge-2 superconductivity for generic singlet states
@@
=-= - $\frac{n\pm 1}{4\pi}(\alpha_{\uparrow}\wedge\mathrm{d}\alpha_{\uparrow}+\alpha_{\downarrow}\wedge\mathrm{d}\alpha_{\downarrow})-\frac{n}{2\pi}\alpha_{\uparrow}\wedge\mathrm{d}\alpha_{\downarrow}$
@@
At this filling, several possibilities to form quantum Hall states exist, which all give rise to superconductivity. One option is for the vortices to form a Laughlin state in each of the valleys, leading to the formation of $2n\pm 1$ decoupled Laughlin states at level 2. This results in charge-2 superconductors which have a large residual topological order ($\det\tilde{K}_{\text{doped}}=16$ in addition to the TO of the superconductor when doping $\nu=2/5$). A perhaps simpler possibility is that vortices spontaneously form $n$ bosonic integer Quantum Hall states with the excess density being accommodated as a level ±$\pm 2$ Laughlin state, with a block-diagonal K-matrix $K_{\text{vortices}}$. This choice generically breaks translations but leads to charge-2 superconductivity for all fillings without additional topological order, thereby naturally matching the superconductor we expect to form on top of the semion crystal. We suspect that in the limit of small electron densities; highly filled bosonic states of vortices become unstable to forming a superfluid and accommodate flux by forming a “vortex lattice” themselves.
@@
$g_{1}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{m+n}\mp\delta)\Phi_{0}$ $+1\rightarrow\mp 1$
$g_{2}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{n+m}\pm\delta)\Phi_{0}$ $+1\rightarrow\pm 1$
@@
$g_{n-1}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{n+m}\mp\delta)\Phi_{0}$ $+1\rightarrow\mp 1$
$g_{n}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{n+m}\pm\frac{\delta}{2})\Phi_{0}$ $+1\rightarrow\pm 2$
@@
results in a charge-2 superconductor. Note that the flux assignment Eq. (66) is consistent with no change in the external flux of $A$ and follows from demanding ×$\nabla\times(a_{\alpha}+a_{\alpha+1})=0$. Using the same arguments as discussed in the main text for $\nu=2/3$, we find that on both sides of the transition, the partons fill an integer number of Landau levels in the doped valley, as summarized in Table 4, where we assumed that translations are broken and doped partons enter in a unique valley. The resulting Lagrangian is
@@
describing the same charge-2 superconductor without leftover topological order as Eq. (35). We remark that the partons $g_{\alpha}$ viewed alone, neglecting the coupling to the spinons, form a $\nu=\mp 2$ bosonic IQH state.
@@ Appendix C Exotic mechanisms for a reentrant integer Quantum anomalous Hall state
Figure 9: Phase transition in the Hubbard-Hofstadter model for $L_{y}=3$. (a) Order parameter $\Delta_{\text{CDW}}$ as a function of the interaction strength $V$ for fixed $U=4t$, $\chi=2048$ and $L_{y}=3$. Inset: Density correlations $C(k_{y}=1,x)=\langle\rho_{-k_{y}}(x)\rho_{k_{y}}(0)\rangle$ for $V=1.4t$ and $V=2.3t$. We show the $k_{y}=1$ correlations, since this corresponds to a cut through the Brillouin zone containing the ordering wave vector $\mathbf{K}$. For the $L_{y}=4$ geometry of the main text, the corresponding cut is $k_{y}=0$. In contrast to the FQH state on the $L_{y}=4$ cylinder discussed in the main text, translations are not even weakly broken in the $(\bar{1}\bar{1}2)$ phase on the $L_{y}=3$ cylinder. (b) The correlation length of the density mode, i.e. of operators with zero charge and spin $(Q,S_{z})=(0,0)$, peaks at a certain interaction strength $V$ which we identify as the critical point (indicated by a dashed line in all panels). The spin correlation length $(Q,S_{z})=(0,\hbar)$ remains finite across the transition. (c) The ratio of $\xi_{2e}$ [$(Q,S_{z})=(2,0)$] to $\xi_{1e}$ [$(Q,S_{z})=(1,\hbar/2)$] correlation length is much larger than the trivially expected $1/2$ and peaks close to the critical point, growing as a function of the bond dimension $\chi$.
[…]
@@ Appendix D Chern numbers upon doping from Diophantine equation
We find that for this flux and filling, $C=-2$ solves the Diophantine equation for any doping $\delta$. We can also show it more explicitly if we write $\delta=1/\ell$ with $\ell\in\mathbb{Z}$, which implies $\phi_{g_{2}}=\frac{2\ell-3}{6\ell}$ while having to fill $62\ell+$ 6 bands. The Chern number after accommodating all the $62\ell+$ 6 electrons is then:
@@
where the first term on the RHS is understood as the modular inverse. The second line follows from $6l=2(2l-3)+(6+2l)$.
mode energy momentum (×$\times 2\pi/L_{x}$) degeneracy
ground state $L_{c}=L_{s}=0$ 00 $1\times(S_{z}=0)$
charge $L_{c}=1$ $-1$ $1\times(S_{z}=0)$
$L_{c}=2$ $-2$ $2\times(S_{z}=0)$
$L_{c}=3$ $-3$ $3\times(S_{z}=0)$
$L_{c}=4$ $-4$ $5\times(S_{z}=0)$
spin $L_{s}=1$ $2$ $3=1\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{s}=2$ $2$ $4=2\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{s}=3$ $3$ $7=3\times(S_{z}=0)+2\times(S_{z}=\pm 1)$
$L_{s}=4$ $4$ $13=5\times(S_{z}=0)+3\times(S_{z}=\pm 1)+1\times(S_{z}=\pm 2)$
mixed $L_{c}=L_{s}=1$ 00 $3=1\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{c}=2,L_{s}=1$ $-1$ $6=2\times(S_{z}=0)+2\times(S_{z}=\pm 1)$
$L_{c}=1,L_{s}=2$ $1$ $4=2\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{c}=L_{s}=2$ 00 $6=4\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
@@
mode energy momentum (×$\times 2\pi/L_{x}$) degeneracy
ground state $L_{c}=L_{s}=1/4$ 00 $2=1\times(S_{z}=\pm 1/2)$
charge $L_{c}=1/4+1$ $-1$ $2=1\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+2$ $-2$ $4=2\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+3$ $-3$ $6=3\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+4$ $-4$ $10=5\times(S_{z}=\pm 1/2)$
spin $L_{s}=1/4+1$ $1$ $2=1\times(S_{z}=\pm 1/2)$
$L_{s}=1/4+2$ $2$ $6=2\times(S_{z}=\pm 1/2)+1\times(S_{z}=\pm 3/2)$
$L_{s}=1/4+3$ $3$ $8=3\times(S_{z}=\pm 1/2)+1\times(S_{z}=\pm 3/2)$
$L_{s}=1/4+4$ $4$ $14=5\times(S_{z}=\pm 1/2)+2\times(S_{z}=\pm 3/2)$
mixed $L_{c}=L_{s}=1/4+1$ 00 $2=1\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+2,L_{s}=1/4+1$ $-1$ $4=2\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+1,L_{s}=1/4+2$ $1$ $6=2\times(S_{z}=\pm 1/2)+1\times(S_{z}=\pm 3/2)$
$L_{c}=L_{s}=1/4+2$ 00 $12=4\times(S_{z}=\pm 1/2)+2\times(S_{z}=\pm 3/2)$

### v2

@@ IV Microscopic lattice model
Next, we study the spin and charge correlation lengths when tuning across the critical point. We find the density correlation length, corresponding to the charge sector $(Q,S_{z})=(0,0)$, to diverge at the critical point as a function of the bond dimension. In contrast, the spin correlation length, $(Q,S_{z})=(0,\hbar)$, remains finite and quickly converges with bond dimension; see Fig. 7(b). This suggests a finite spin gap across the transition, consistent with the continuous transition discussed in Sec. II. Furthermore, the ratio between the charge $(Q,S_{z})=(2,0)$ and charge $(Q,S_{z})=(1,\hbar/2)$ correlation length peaks at the critical point, growing with bond dimension and are, in particular, larger than the trivial expectation of 1/2, which arises when the charge-2 correlations would be only inherited from charge-1 correlations; Fig. 7(c). This behavior is expected when charge-2 excitations become energetically favorable compared to spinful charge-1 excitations. We emphasize that the large spin gap in both the FCI and the semion crystal protects these phases against perturbations. Accordingly, we expect them to be stable beyond the $\mathrm{SU(2)}$ symmetric model Eq. (51) considered here. Since the spin gap remains finite across the transition, we anticipate this stability to be inherited upon doping in the vicinity of the critical point.
@@
We note that even in the strict limit of spinless fermions, the system could undergo the same FCI to semion crystal transition, enabling an identical anyon-superconductivity mechanism in its vicinity, just as in the spinful case. The relevant energy scale near the transition is set by a finite charge $Q=1e$ gap. Finding a microscopic model that stabilizes such spinless semion crystal and perhaps realizes a continuous transition from a 2/3 FCI phase, remains an exciting open challenge.
@@ V Experimental implications
Therefore, from a coarse-grained point of view, both our setting and MoTe2 host the fractional $2/3$ state, a proximate superconductor on doping, and nearby insulators without Hall conductance. However, some important differences between these systems should be noted, even at the level of symmetries. We have assumed spin rotation symmetry, while the valley polarized MoTe2 has stongly-broken spin rotation symmetry due to spin valley locking. Can the two be related? The first observation is that since our model has a spin gap, weak breaking of spin-rotation symmetry leaves the phases intact. For instance, the topological order of the chiral spin liquid is stable when spin rotation symmetry is broken [34], despite the fact that spin transport is not quantized any longer. Next, we note that for the strong spin-rotation symmetry breaking relevant to tMoTe2, our mechanism provides a general guideline: if the 2/3 FCI is in proximity to a specific charge density wave - the semion crystal - then the charge 2e/3 anyon becomes energetically cheaper than a pair of separated e/3 quasiparticles. Doping such an FCI should favor superconductivity. Repulsive interactions are indeed expected to drive FCIs towards CDW order, the key is to realize the correct type of CDW, (see Appendix I). A promising strategy is to search the phase diagram of MoTe2 for insulating states that continue to break time reversal symmetry but exhibit a vanishing Hall response, which can be a potential candidate for this exotic CDW and examine if we have a continuous (or weakly first order) transition into the 2/3 FCI. Indeed, it has already been speculated that precisely such exotic CDWs could arise in MoTe2 [71]. Our framework links that putative CDW directly to enhanced Cooper pairing in the neighboring FCI, providing a concrete route to accounting for superconductivity.
@@
As an alternative scenario, superconductivity in this system could instead evolve from the intervening chiral metal [86, 19]. To distinguish these scenarios, it will be helpful to have more information on the energy ordering of quasiparticles. A simple (but indirect and imperfect) way to do this is via the observation (or absence) of daughter FCI plateaus, as the hierarchical sequence changes depending on which anyon is lowest energy. If the minimal charge 1/3 anyon is the cheapest, then one obtains the usual Haldane-Halperin hierarchy [23], and several of these are observed for $\nu<2/3$. This is consistent with that being the lowest charge excitation on the electron doping (smaller $\nu$) side. However on the opposite side, incompressible states at fillings predicted by the 2/3 anyon hierarchy, such as $\nu=4/5$ would provide strong evidence for 2/3 anyons to be cheaper for hole doping (see Appendix B), regardless of the mechanism. This energy ordering would be consistent with the observation of superconductivity for $\nu>2/3$. Other potential signatures of the energy ordering of anyons could come from spectroscopic probes [28, 54, 55, 27, 52, 56] or by measuring the electric charge via shot-noise measurements [8, 61] or via single electron transitor measurements [47, 78].
@@
In the MoTe2 experiments, a re-entrant integer QAH plateau (RIQAH) was reported [86] between the 2/3 and 3/5 FQAH states. Although additional confirmation is warranted—particularly since edge-equilibration effects could influence the measurement—taking the observation at face value, this state could potentially arise from the usual plateau transition driven by 1/3 anyons. In our formalism, a different possibility is suggested. Stacking a Chern-1 QAH layer on top of the 2/5 to semion-crystal transition also reproduces such a feature (see Appendix C). The stacking converts the 2/5 into a 3/5 FQAH state, while the CDW, after stacking, reproduces the Hall conductance of the RIQAH. Note however, the presence of a background topological order in this scenario implies that this phase is more appropriately designated RIQAH.
@@ VI Outlook & Discussions
One of the biggest surprises of our work is the numerical discovery of the semion crystal in simple Hubbard models at fractional filling. One may view this phase as a period-three electron CDW on which the residual spins form a chiral spin liquid, akin to the cluster-Mott physics proposed for 1T-TaS2 [39]. Finding a stable semion crystal in such a minimal model highlights it as a viable state even competing with conventional crystals and FQH states. Elucidating the properties and stability of such semion crystals at other fillings or in the absence of a moiré potential (i.e. the formation of a Wigner crystal supporting a CSL) is an interesting direction.
@@
Acknowledgments.— We thank J. Dong, P.J. Ledwith, J. Feldmeier, A. Imamoglu, G. Shavit, J. May-Mann, T. Helbig, and J. Alicea for insightful discussions. A.V. would like to thank Mike Zaletel, Stefan Divic, Zhaoyu Han, Darius Shi, Carolyn Zhang and Yahui Zhang for discussions and collaborations on related topics. C.K. acknowledges funding from the Swiss National Science Foundation (Postdoc.Mobility Grant No. 217884). A.V. is supported by a Simons Investigator award, the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651440, A.V.). F.P. and M.K. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, TRR 360 – 492547816 and DFG grants No. KN1254/1-2, KN1254/2-1, the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 851161), the European Union (grant agreement No 101169765), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@
Appendix A Charge-2e superconductivity for generic singlet states
@@
$=$- $\frac{n\pm 1}{4\pi}(\alpha_{\uparrow}\wedge\mathrm{d}\alpha_{\uparrow}+\alpha_{\downarrow}\wedge\mathrm{d}\alpha_{\downarrow})-\frac{n}{2\pi}\alpha_{\uparrow}\wedge\mathrm{d}\alpha_{\downarrow}$
@@
At this filling, several possibilities to form quantum Hall states exist, which all give rise to superconductivity. One option is for the vortices to form a Laughlin state in each of the valleys, leading to the formation of $2n\pm 1$ decoupled Laughlin states at level 2. This results in charge-2e superconductors which have a large residual topological order ($\det\tilde{K}_{\text{doped}}=16$ in addition to the TO of the superconductor when doping $\nu=2/5$). A perhaps simpler possibility is that vortices spontaneously form $n$ bosonic integer Quantum Hall states with the excess density being accommodated as a level $\pm 2$ Laughlin state, with a block-diagonal K-matrix $K_{\text{vortices}}$. This choice generically breaks translations but leads to charge-2e superconductivity for all fillings without additional topological order, thereby naturally matching the superconductor we expect to form on top of the semion crystal. We suspect that in the limit of small electron densities; highly filled bosonic states of vortices become unstable to forming a superfluid and accommodate flux by forming a “vortex lattice” themselves.
@@
$g_{1}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{m+n}\mp\delta)\Phi_{0}$ $1\rightarrow\mp 1$
$g_{2}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{n+m}\pm\delta)\Phi_{0}$ $1\rightarrow\pm 1$
@@
$g_{n-1}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{n+m}\mp\delta)\Phi_{0}$ $1\rightarrow\mp 1$
$g_{n}$ $\frac{1}{m+n}+\delta$ $(\frac{1}{n+m}\pm\frac{\delta}{2})\Phi_{0}$ $1\rightarrow\pm 2$
@@
results in a charge-2e superconductor. Note that the flux assignment Eq. (66) is consistent with no change in the external flux of $A$ and follows from demanding $\nabla\times(a_{\alpha}+a_{\alpha+1})=0$. Using the same arguments as discussed in the main text for $\nu=2/3$, we find that on both sides of the transition, the partons fill an integer number of Landau levels in the doped valley, as summarized in Table 4, where we assumed that translations are broken and doped partons enter in a unique valley. The resulting Lagrangian is
@@
describing the same charge-2e superconductor without leftover topological order as Eq. (35). We remark that the partons $g_{\alpha}$ viewed alone, neglecting the coupling to the spinons, form a $\nu=\mp 2$ bosonic IQH state.
@@ Appendix C Exotic mechanisms for a reentrant integer Quantum anomalous Hall state
Figure 9: Phase transition in the Hubbard-Hofstadter model for $L_{y}=3$. (a) Order parameter $\Delta_{\text{CDW}}$ as a function of the interaction strength $V$ for fixed $U=4t$, $\chi=2048$ and $L_{y}=3$. Inset: Density correlations $C(k_{y}=1,x)=\langle\rho_{-k_{y}}(x)\rho_{k_{y}}(0)\rangle$ for $V=1.4t$ and $V=2.3t$. We show the $k_{y}=1$ correlations, since this corresponds to a cut through the Brillouin zone containing the ordering wave vector $\mathbf{K}$. For the $L_{y}=4$ geometry of the main text, the corresponding cut is $k_{y}=0$. In contrast to the FQH state on the $L_{y}=4$ cylinder discussed in the main text, translations are not even weakly broken in the $(\bar{1}\bar{1}2)$ phase on the $L_{y}=3$ cylinder. (b) The correlation length of the density mode, i.e. of operators with zero charge and spin $(Q,S_{z})=(0,0)$, peaks at a certain interaction strength $V$ which we identify as the critical point (indicated by a dashed line in all panels). The spin correlation length $(Q,S_{z})=(0,\hbar)$ remains finite across the transition. (c) The ratio of $\xi_{2e}$ $(Q,S_{z})=(2,0)$] to $\xi_{1e}$ $(Q,S_{z})=(1,\hbar/2)$] correlation length is much larger than the trivially expected $1/2$ and peaks close to the critical point, growing as a function of the bond dimension $\chi$.
Recent experiments have observed signatures consistent with a Reentrant Integer Quantum Anomalous Hall effect (RIQAH), appearing between $\nu=3/5$ and $\nu=2/3$. If one considers the $\nu=2/3$ state as a FQH state of holes forming out of a fully occupied band $\nu=1-1/3$, it is natural to describe the transition $\nu=2/3\leftrightarrow\text{RIQAH}$ as a crystallization transition $\nu=1-1/3\leftrightarrow\nu=1+\text{CDW}$ for which the $Q=e/3$ anyon is active [71]. An even more exotic option can arise by approaching the RIQAH state from $\nu=3/5$. This transition could take place if a $\nu=-2/5$ state transitions to a topological crystal as outlined in the main text, leaving behind an IQH response $\text{IQH}-2/5\leftrightarrow\text{IQH+semion crystal}$. In such a case, the Q=2e/5Q=-2e/5 anyon could be the lowest-lying excitation, thereby presenting a pathway to a superconducting state near the RIQAH phase.
@@ Appendix D Chern numbers upon doping from Diophantine equation
We find that for this flux and filling, C=2C=-2 solves the Diophantine equation for any doping $\delta$. We can also show it more explicitly if we write $\delta=1/\ell$ with $\ell\in\mathbb{Z}$, which implies $\phi_{g_{2}}=\frac{2\ell-3}{6\ell}$ while having to fill $2\ell+6$ bands. The Chern number after accommodating all the $2\ell+6$ electrons is then:
@@
where the first term on the RHS is understood as the modular inverse. The second line follows from 6l=2(2l3)+(6+2l)6l=2(2l-3)+(6+2l). We remark that simple solutions to the Diophantine equations resulting in superconductors are found only for specific choices of electron density $n_{e}$ and background-flux $n_{\Phi}$ per unit-cell. Consequently, we expect that the superconductor has to accommodate excess magnetic field by forming vortices.
mode energy momentum ($\times 2\pi/L_{x}$) degeneracy
ground state $L_{c}=L_{s}=0$ 0 $1\times(S_{z}=0)$
charge $L_{c}=1$ 1-1 $1\times(S_{z}=0)$
$L_{c}=2$ 2-2 $2\times(S_{z}=0)$
$L_{c}=3$ 3-3 $3\times(S_{z}=0)$
$L_{c}=4$ 4-4 $5\times(S_{z}=0)$
spin $L_{s}=1$ 22 $3=1\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{s}=2$ 22 $4=2\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{s}=3$ 33 $7=3\times(S_{z}=0)+2\times(S_{z}=\pm 1)$
$L_{s}=4$ 44 $13=5\times(S_{z}=0)+3\times(S_{z}=\pm 1)+1\times(S_{z}=\pm 2)$
mixed $L_{c}=L_{s}=1$ 0 $3=1\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{c}=2,L_{s}=1$ 1-1 $6=2\times(S_{z}=0)+2\times(S_{z}=\pm 1)$
$L_{c}=1,L_{s}=2$ 11 $4=2\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
$L_{c}=L_{s}=2$ 0 $6=4\times(S_{z}=0)+1\times(S_{z}=\pm 1)$
@@
mode energy momentum ($\times 2\pi/L_{x}$) degeneracy
ground state $L_{c}=L_{s}=1/4$ 0 $2=1\times(S_{z}=\pm 1/2)$
charge $L_{c}=1/4+1$ 1-1 $2=1\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+2$ 2-2 $4=2\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+3$ 3-3 $6=3\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+4$ 4-4 $10=5\times(S_{z}=\pm 1/2)$
spin $L_{s}=1/4+1$ 11 $2=1\times(S_{z}=\pm 1/2)$
$L_{s}=1/4+2$ 22 $6=2\times(S_{z}=\pm 1/2)+1\times(S_{z}=\pm 3/2)$
$L_{s}=1/4+3$ 33 $8=3\times(S_{z}=\pm 1/2)+1\times(S_{z}=\pm 3/2)$
$L_{s}=1/4+4$ 44 $14=5\times(S_{z}=\pm 1/2)+2\times(S_{z}=\pm 3/2)$
mixed $L_{c}=L_{s}=1/4+1$ 0 $2=1\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+2,L_{s}=1/4+1$ 1-1 $4=2\times(S_{z}=\pm 1/2)$
$L_{c}=1/4+1,L_{s}=1/4+2$ 11 $6=2\times(S_{z}=\pm 1/2)+1\times(S_{z}=\pm 3/2)$
$L_{c}=L_{s}=1/4+2$ 0 $12=4\times(S_{z}=\pm 1/2)+2\times(S_{z}=\pm 3/2)$